Sue's father is 6 times as old as she is. In 4 years, he will be 4 times as old. How old are they now?
Answer by TimothyLamb(4379)

f = 6s
f + 4 = 4(s + 4)
f + 4 = 4s + 16
f = 4s + 12
---
6s = 4s + 12
2s = 12
---
Answer:
s = 6
f = 36
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php